OpenStudy (anonymous):

2(h – 8) – h = h – 16 h = 8 h = –8 h = infinitely many solutions h = no solution would the answer be -8?

OpenStudy (anonymous):

2(8-8)-8=8-16 2(0)-8=-8 -8=-8

OpenStudy (anonymous):

the variable is gone

OpenStudy (anonymous):

so it would be no solution?

OpenStudy (anonymous):

would it?

OpenStudy (anonymous):

can you explain how you got your answer,please.

OpenStudy (anonymous):

how did i get the answer? i got the answer by plugining 8 into the equation

OpenStudy (anonymous):

alright thanks.